OpenStudy (anonymous):

Soybean is 12% protein; cornmeal is 6% protein. how many pounds of each should be mixed together in order to get 240-lb mixture that is 8% protein?

OpenStudy (anonymous):

two equations (1) 240=s+c (2)19.2=.12s+.06c solve

OpenStudy (amistre64):

cornmeal@6% = 240(.08-.12)/(.06-.12)

OpenStudy (anonymous):

amistre!!! help me with my integral

OpenStudy (anonymous):

Please :D

OpenStudy (anonymous):

Should i solve that substitution?

OpenStudy (anonymous):

Can you explain this some more, please

OpenStudy (amistre64):

ideally, youd solve the system of equations by substitution yes a+b=t; a=t-b ax +by=tz (t-a)x + by = tz tz -bx +by = tz b(y-x) = t(z-x) b@y = t(z-x)/(y-x) cornmeal@6% = 240(.08-.12)/(.06-.12)

OpenStudy (amistre64):

t-a means t-b ... typoed it :)

OpenStudy (anonymous):

okay so i would, multiply 240 by (.08-.12) and then divide by (.06-.12)... or divide both (.08-.12)/(.06-.12) first and then multiply by 240?

OpenStudy (amistre64):

its the same regardless of how you go at it. You can try it out both ways and check that the answers are the same. 240 (.08-.12) 240(.08-.12) -------- * (.08-.12) = --------- * 240 = ------------ etc (.06-.12) (.06-.12) (.06-.12)